118. 3+ 1/x-2 = 2x-3/x-2 what do i do with this equation and how do i work it i do not understand can you please help??? :) thank you for your hard work

Answer by stanbon(57395)

3+ 1/(x-2) = (2x-3)/(x-2)
Multiply thru by (x-2) to get rid of the denominators:
3(x-2)+1 = 2x-3
Simplify:
3x-6+1 = 2x-3
Subtrac 2x from both sides:
x-5=-3
x=2
Cheers,
Stan H.